Calculate the equivalent weight of an acid, if 0.275 grams of the acid are completely titrated with 27.8 ml of 0.214 N NaOH. Show all work
Answer #1

Number of equivalents of acid = number of equivalents of NaOH =Normality * Volume = 0.214 * 27.8 * 10-3 = 5.949 * 10-3

Equivalents of acid = mass of acid / equivalent weight

5.949 * 10-3 = 0.275 / Equivalent weight

Equivalnt weight = 46.226 g/eq